# Surg-3M: A Dataset and Foundation Model for Perception in Surgical Settings

### v2

@@ 6.1 Ablation Studies
SurgFM design choices. We performed incremental ablation studies that test the choice of backbone (ViT vs ConvNeXt), self-supervision method (DINO vs DINOv2 vs MAE), and our proposed augmented distillation (Sec. 4). The results are shown in Table 2. DINO outperformed DINOv2 and MAE in all the downstream tasks and metrics evaluated, with improvements $>$8 percentage points (pp) in mAP in both tool detection and action recognition. ConvNeXt-L achieved comparable results to ViT-L in phase recognition, and improved $>$2pp (F1-score) in surgical tool detection and $>$1pp (mAP) in surgical action recognition. Our proposed augmentation achieved comparable results in phase recognition, and improved $>$1pp (mAP) in tool detection and $>$1pp (F1-score) in action recognition.
Pretrained weights. We compare the effects of employing pretrained weights from ImageNet in the ConvNeXt-L backbone with those of using Surg-3M. As shown in Tables 3, 4, and 5, SurgFM (which also has a ConvNeXt-L backbone, but is pretrained on Surg-3M) outperformed the ConvNeXt-L backbone pretrained on ImageNet in all downstream tasks and datasets. This finding suggests that SurgFM serves as a superior starting point for surgery-specific downstream task fine-tuning compared to models pretrained on general-purpose datasets like ImageNet. This is clearly shown in the surgical tool presence detection experiments (Table 4) on the GraSP dataset, where SurgFM outperformed a ConvNeXt-L pretrained on ImageNet by more than 4pp in mAP.
6.2 Comparison with SotA
The performance of our SurgFM compared to the state of the art in each downstream task and dataset is shown in Fig. 1 and Tables 3, 4, and 5.
Table 5: Performance evaluation on surgical action recognition. Two datasets are used: CholecT50 [42] and SAR-RARP50 [48]. The metrics are computed as described in Sec. 5.1. The evaluation protocol and the meanings of G, S, and F are explained in Sec. 5.2. The superscript denotes that the SurgFM backbone is frozen during fine-tuning. The subscript ± indicates that both past and future frames are used [48].
Method Type CholecT50 SAR-RARP50
mAP (%) Accuracy (%) F1@10 (%)
ConvNeXt-L± [36] G $62.9\pm 2.3$ $71.9$ $70.4$
Tripnet [43] S 58.8$\pm$3.1 - -
Rendezvous [44] $60.4\pm 2.8$
Attention Tripnet [44] $61.1\pm 2.0$
SummerLab-AI± [48] - $\mathbf{81.5}$ $\mathbf{84.1}$
Uniandes± [48] - $78.6$ $82.3$
MS-TCN± [48, 16] - $69.0$ $70.7$
LoViT± [48, 35] - $59.8$ $43.0$
SurgFM (ours) F $\mathbf{64.2\pm 1.7}$ $70.2$ $68.4$
${}_{\pm}^{*}$ (ours) $63.7\pm 1.7$ $72.4$ $74.3$
SurgFM± (ours) $\mathbf{64.2\pm 1.7}$ $74.1$ $74.3$
Surgical phase recognition. As shown in Table 3, SurgFM outperformed all specialist [54, 27, 52, 26, 62, 12, 20, 29, 30, 35] and foundation [49, 24, 5] models in Cholec80 ($>$3pp Jacc.), AutoLaparo ($>$8pp Jacc.), and M2CAI16 ($>$4pp Jacc.). These results demonstrate that SurgFM is able to effectively extract features that are useful for different surgical procedures, capitalizing on the diversity of surgery types present in the Surg-3M dataset.
In GraSP, SurgFM demonstrated performance comparable (with an mAP difference of $<$1pp) to that of the methods presented in [17, 55, 4]. However, these competing methods [17, 55, 4] utilize both past and future frames (offline), whereas our SurgFM is restricted to observing only the current and preceding frames, reflecting its real-time operational focus, and its broader range of application [35]. Despite being constrained by its limited temporal perspective, SurgFM achieves comparable performance with these more information-rich approaches.
Surgical tool presence detection. As shown in Table 4, SurgFM outperformed all specialist [54, 27, 44] and foundation [49] models in Cholec80, and achieved a comparable performance to [44, 4] in GraSP.
Surgical action recognition. As shown in Table 5, SurgFM outperformed all specialist model [43, 44] in CholecT50 ($>$3pp mAP), and achieved a competitive performance to [48] in SAR-RARP50 ($<$7pp difference in accuracy).
@@ 6.2 Comparison with SotA
Table 6: Leaderboard for the proposed tasks on Surg-3M. Multi-label (35 classes) video classification of procedure types and binary classification of surgery types.
@@
mAP (%) F1 (%) Accuracy (%) F1 (%)
SlowFast [17] 22.0 23.9 88.5 87.5
TimeSformer [6] 42.1 37.5 93.2 92.7
MViTv2 [34] 49.5 41.8 95.8 94.6
Video Swin Transformer [37] 51.4 47.9 98.8 98.7
SurgFM-Vid (ours) $\mathbf{57.8}$ $\mathbf{49.3}$ $\mathbf{98.9}$ $\mathbf{98.9}$
Frozen SurgFM model. As shown in Tables 3, 4, and 5, the frozen version of SurgFM, denoted SurgFM, is indeed a foundation model. It works on all datasets and tasks without retraining. In Cholec80, SurgFM outperforms specialist SotA architectures such as EndoNet, PhaseNet, and SV-RCNet, OHFM, TeCNO. In AutoLaparo, SurgFM outperforms all the six specialist models we compared it with. In M2CAI16, SurgFM outperforms two specialist SotA models, PhaseNet and SV-RCNet. In SAR-RARP50, SurgFM outperforms MS-TCN and LoViT.
Data scaling. We investigated the extent to which annotation effort can be alleviated by leveraging our pretrained SurgFM model. To do so, we fine-tuned it on downstream tasks with 10%, 50%, and 100% of the annotated data (Fig. 5). With 50% of the data, the performance decreased a maximum of 4.4pp across all datasets and tasks except GraSP, which has a total of four videos for training at 100%, and where performance dropped by 9.6pp. The closest result was achieved in phase recognition (Cholec80), with SurgFM achieving an accuracy of 92.3% on 50% of the data, only 0.1pp below prior SotA with 100% of the data [35]. Also at 50%, our SurgFM outperformed prior SotA [35] in AutoLaparo at 100%. These results suggest that the features learned by SurgFM from Surg-3M are highly informative. Comparing our results with other foundation models reported in the literature, SurgFM achieved $>$6pp (F1-score) than Ramesh et al. [49] in phase recognition (Cholec80), despite using only 10% of the data compared to their 12.5%. Furthermore, SurgFM also outperformed Ramesh et al. by $>$6pp (mAP) on tool presence detection (Cholec80) at 10%.
@@ 6.3 Leaderboard
To establish a baseline for the novel tasks presented in Sec. 3, we conducted an evaluation of SotA approaches on Surg-3M (Table 6), which serves as a benchmark for future research endeavors. By leveraging the robust representations extracted from SurgFM, our video classification model, SurgFM-Vid (Sec. 4), demonstrated significant improvements ($>$6 mAP) over all the leading video classification models tested on the multi-label procedure type classification task. On the binary surgery type classification task, our approach achieved comparable accuracy to the top performer, the Video Swin Transformer [37]. The incorporation of frame-level features extracted by SurgFM, in conjunction with the strategic application of prior knowledge through weighted feature embeddings informed by frame typicality, leads to substantial performance gains, underscoring the value of SurgFM as a critical component in surgical video analysis tasks. Despite these improvements, our results also highlight the challenges associated with surgical procedure classification, underscoring the need for continued innovation in this area.
@@ 7 Conclusion
Figure 5: Data scaling. Results of SurgFM when fine-tuned on the downstream tasks with different percentages of annotated training samples. The radius of the circles, computed following $2\log(\sigma+1)$, represents the standard deviation ($\sigma$) of each metric estimated over five cross-validation folds. Prior SotA is computed using 100% of the available data. Best viewed online.
Building on a large corpus of publicly available surgical videos, this work has two main contributions: we have created a large-scale and diverse surgical video dataset, Surg-3M, and a foundation model, SurgFM. Our SurgFM outperformed specialist architectures on four datasets and three downstream tasks, demonstrating its potential as a valuable tool for computer vision research in surgery. Furthermore, our experimental results show that SurgFM can maintain its performance with significantly reduced data annotation needs (50% or even 10% for some tasks), highlighting the efficiency of our approach. This work showcases the value of Surg-3M and SurgFM as a foundation for future research in surgical computer vision. To further improve these contributions, we plan to address key limitations of our current approach, including adapting SurgFM to pixel-wise tasks and enhancing the self-distillation approach to better adapt to surgical data. Ultimately, this work highlights the importance of publicly available surgical videos in developing more effective models for computer vision research in surgery, and we believe that Surg-3M and SurgFM will serve as catalysts for future advancements in this field.
@@
Your expertise includes identifying and categorizing surgical interventions based on clinical
descriptions and procedural contexts.
Here is a list of 35 possible surgical procedure types: pancreatectomy, pancreaticoduodenectomy,
splenectomy, ampullectomy, hepatectomy, nephrectomy, low anterior resection, colectomy,
abdominoperineal resection, pulmonary lobectomy, hartmanns, prostatectomy, gastric bypass,
duodenal switch, gastrectomy, small bowel resection, hernia repair, ulcer repair, cholecystectomy,
appendectomy, ileocolic resection, cecectomy, myomectomy, hysterectomy, nissen fundoplication,
adrenalectomy, thymectomy, rectopexy, adhesiolysis, esophagectomy, cystectomy, jejunostomy, ileorectal
anastomosis, kidney transplant, vaginectomy.
Based on the description of the surgical video: <video title>, determine the most likely procedure type
from the list. Focus on matching the description to the procedure type that best aligns with the
terminology and context provided.
Figure S1: ChatGPT prompt employed to match video titles to procedure types. The title of the video to be matched is inserted where the $<$video title$>$ tag is located.
@@ Appendix A Surgical Background
Appendix B Details of Data Curation
This section provides details of the implementation of our data curation pipeline.
@@ B.1 Video Classification
Video summarization. To improve the efficiency of annotating surgical and non-surgical videos, we obtained 4$\times$4-image video storyboards (i.e. a single image containing a collage of key video frames) for all the collected videos using the method described in [18]. Storyboards enabled us to quickly determine whether a video contained substantial surgical footage, thereby avoiding the need for complex analysis on the entire sequence.
@@
Performance of video storyboard classification models. We trained five video storyboard classification models (different data splits) to categorize the videos as either surgical or non-surgical. Each fold was split into training, validation, and testing sets with ratios of 0.8, 0.1, and 0.1, respectively. The average F1-score of the video storyboard classification models was $94.75\%\pm 1.1$. The results for each cross-validation fold are shown in Table S1.
Table S1: Performance of the video storyboard classification models across five folds.
@@ B.2 Video Selection and Trimming
Performance of frame classification models. We trained five frame classification models to classify a video frame as either surgical or non-surgical. Each model of the five was trained in a different training-validation-testing split of the data, with a split rate of 0.8, 0.1, and 0.1. As shown in Table S2, the average F1-score of the frame classification models was $95.64\%\pm 0.94$.
Table S2: Performance of frame classification models across five folds.
@@
B.3 Video Annotation
For the surgery type, a video is considered to be robotic if the video title includes any of the following keywords: Robotic, Robot, Robo, Hugo, Versius, Senhance, Telerobotic, Console, and da Vinci. The search for these terms in the video titles was case-insensitive. The remaining videos, with titles that do not include any of these keywords, are manually verified to ensure that they are manual surgical procedures. For the surgical procedure type, we cross-reference the video titles with the predefined list of procedures. When exact title matches are not found, we leverage the capabilities of the ChatGPTv4 API to perform a more nuanced analysis, incorporating a customized prompt as shown in Fig. S1. All annotations are manually quality controlled.
After curation, the final Surg-3M dataset resulted in $4194$ surgical videos, 94% of which are $1280\times 720p$, the remaining $6\%$ have varied resolutions, with a minimum of $640\times 480$ pixels.
Appendix C SurgFM Pretraining: Implementation Details
This section provides an overview of the pretraining process for the SurgFM model, and the image preprocessing pipeline employed.
SurgFM pretraining. For the pretraining of SurgFM, we utilized eight NVIDIA V100 GPUs (32 GB) with a batch size of 24 per GPU. We employed AdamW as the optimizer with a teacher temperature of 0.04, fp16 precision, an initial learning rate of 5e-4 (after warm-up), and a minimum learning rate of 1e-6. The model was trained for 60 epochs, including 10 warm-up epochs. This number of epochs enables the model to converge and stabilize its training loss on our dataset and pretext task. The model with the lowest training loss was selected as the final model.
Figure S2: Image preprocessing in SurgFM. We automatically detect and crop out the user interface information from the surgical frames.
Image preprocessing. We utilize our trained non-surgical bounding-box detection model to identify and obliterate UI elements, as demonstrated in Fig. S2.
Performance of non-surgical content detection models. We trained five non-surgical content detection models to detect and obliterate non-surgical regions in surgical video frames. Each model was trained on a different training-validation-testing split of the data, with a split ratio of 0.8, 0.1, and 0.1, respectively. The average mAP50 of the five models was $79.29\%\pm 3.4$, and the average mAP50-95 was $66.18\%\pm 2.98$. The results for all the cross-validation folds are shown in Table S3.
Table S3: Performance of bounding box detection models across five folds.
Fold Images Instances Box Precision (%) Box Recall (%) mAP50 (%) mAP50-95 (%)
@@ Appendix C SurgFM Pretraining: Implementation Details
Appendix D Downstream Tasks: Data Splits and Fine-tuning
In this section, we provide the details of the datasets used to evaluate each downstream task along with their corresponding data splits.
Surgical phase recognition. For M2CAI16, we follow the data splits in [26, 54, 30], dividing the dataset into 27 training videos and 14 testing videos, respectively. For Cholec80, we adopt the data splits in [54, 12, 35, 27], allocating 40 videos for training, eight for validation, and 32 for testing. For AutoLaparo, we follow the standard split of ten training videos, four for validation, and seven for testing [57]. For GraSP, we adopt the data splits specified in [4], with four videos for training, four for validation, and five for testing.
Surgical tool presence detection. The data splits for Cholec80 and GraSP are identical to those employed in the surgical phase recognition task.
Surgical action recognition. For CholecT50, we follow the standard 5-fold data splits as proposed in [42]. For SAR-RARP50 [48] we use its recommended split of 27 videos for training, 13 for validation, and 10 for testing.

### v3

@@ 5.3 Ablation Studies
We conducted incremental ablation studies on the surgical phase recognition (AutoLaparo) and surgical semantic segmentation (CholecSeg8k) tasks (see Table 5).
Table 5: Ablation studies. We froze the backbone and only fine-tuned the task-specific heads (i.e., a linear head for phase recognition and a FPN head for semantic segmentation). The column “Curation” marks whether the raw surgical videos were processed with the pipeline in Fig. 2. The column “Aug” shows whether the augmented distillation method in Fig. 4 was applied.
Dataset Curation Aug Backbone AutoLaparo CholecSeg8k
Acc/F1 mDice
IN-1K N/A ConvNeXt-L 63.6/53.063.6/53.063.6 / 53.0 64.464.464.4
LEMON ConvNeXt-L 71.7/61.471.7/61.471.7 / 61.4 67.467.467.4
LEMON ConvNeXt-L 75.3/65.975.3/65.975.3 / 65.9 68.768.768.7
LEMON ViT-L 75.6/66.175.6/66.175.6 / 66.1 61.261.261.2
LEMON ConvNext-B 74.7/64.574.7/64.574.7 / 64.5 68.568.568.5
LEMON ConvNeXt-L $\mathbf{76.4/66.9}$ $\mathbf{71.9}$
The significance of LEMON. We compare the effects of using LEMON for pretraining the surgical foundation model against those obtained with ImageNet-1K [13] (row 1 and row 3 in Table 5). The model trained on LEMON outperformed the one pretrained on ImageNet by a large margin (+12.9pp F1-score on phase recognition, +4.3pp mDice on semantic segmentation). This finding suggests that pretraining on LEMON offers a superior approach for surgery-specific foundation model training compared to pretraining on general-purpose datasets like ImageNet.
The importance of rigorous data curation. We analyze the importance of our data curation pipeline by comparing LemonFM against a model pretrained on an uncurated version of LEMON (i.e., the raw surgical YouTube videos without our data curation). Our data curation pipeline (Fig. 2) yields substantial improvements. Specifically, the curated dataset leads to an increase of 4.5pp F1-score in the phase recognition task and 1.3pp of mDice in the semantic segmentation task (rows 2 and 3 in Table 5).
The effectiveness of the augmented distillation. We compared the effectiveness of employing our proposed augmented distillation method (Fig. 4) against a vanilla DINO [8]. LemonFM consistently outperformed the vanilla DINO baseline, with the largest gain (+3.2pp mDice) in semantic segmentation (rows 3 and 6 in Table 5).
LemonFM design choices. We performed ablation studies to evaluate the effects of two design choices for surgical foundation model pretraining: backbone architecture, ViT vs. ConvNeXt, and self-supervision method, generative-based [21] vs. discriminative-based [8]. As shown in row 4 and 6 of Table 5, when pretrained on LEMON with our augmented knowledge distillation, the ConvNeXt backbone consistently outperformed the ViT-based counterpart. The improvement in performance is particularly notable in the pixel-wise semantic segmentation task (+10.7pp mDice). In Tables 2, 3, and 4, we compare methods whose pretext task is generative-based (MAE and EndoViT) against methods whose pretext task is discriminative-based (classification) such as Endo-FM, SurgeNetXL and our LemonFM. The results show that discriminative‑based models yield a substantial performance advantage, especially when the backbone is frozen, indicating generative objectives such as MAE are less effective at capturing surgical visual cues.
@@
Table 6: Leaderboard for the proposed tasks on LEMON. Multi-label (35 classes) video classification of procedure types and binary classification of surgery types.
@@
mAP/F1 Acc/F1
SlowFast 22.0 /23.9 88.5 /87.5
TimeSformer 42.1/37.5 93.2 /92.7
MViTv2 49.5/41.8 95.8 /94.6
Video Swin Transformer 51.4/47.9 98.8/98.7
LemonFM-Vid $\mathbf{57.8/49.3}$ $\mathbf{98.9/98.9}$
@@ 5.4 Leaderboard
We benchmarked state-of-the-art video classification models on LEMON (Table 6), including SlowFast [15], TimeSformer [6], MViTv2 [34] and Video Swin Transformer [37]. Leveraging the robust representations extracted from LemonFM, our video classification model, LemonFM-Vid (Sec. 3), improved multi‑label procedure classification by 6.4pp mAP over all leading models and matched the top performance of the Video Swin Transformer on the binary surgery type classification task. The incorporation of frame-level features extracted by LemonFM, in conjunction with the weighted feature embeddings informed by frame typicality, led to substantial performance gains, underscoring the value of LemonFM for surgical video analysis tasks. Nonetheless, the mAP score of 57.8% achieved by LemonFM-Vid (the top performer method) highlights the challenges associated with our newly proposed procedure type video classification task, underscoring the need for continued innovation in this area.
@@ 6 Conclusion
Building on a large corpus of publicly available surgical videos, in this work we presented three main contributions: 1) we proposed a novel data curation pipeline to curate online surgical videos; 2) leveraging this pipeline, we created a large-scale and diverse surgical dataset, LEMON; and 3) we developed a foundation model specific for surgical applications, LemonFM. Our results show that LemonFM outperformed both state-of-the-art surgical foundation models and task-specific architectures across diverse surgical downstream tasks. Even when fine-tuned with only half of the labeled data, LemonFM maintained its advantages. These findings demonstrate LemonFM’s effectiveness and verify the quality of LEMON as a valuable resource for computer vision research in surgery. We expect that both LEMON and LemonFM will serve as a foundation for future research in surgical computer vision and facilitate the advancement of autonomous robotic surgery systems. Future work will concentrate on the development of a surgery-specific video foundation model.
@@
Figure 1: Diversity and procedure prevalence in LEMON. Representative samples from various procedures, demonstrating the diverse range of cases in our curated dataset (left, right). Distribution of surgical frames by procedure type (center).
@@ Appendix A Surgical Background
Appendix B Data Curation Details
This section provides details of our data curation pipeline and presents LEMON’s procedure diversity and prevalence (Fig. 1).
@@ B.1 Video Classification
Video summarization. To improve the efficiency of annotating surgical and non-surgical videos, we obtained 4$\times$4-image video storyboards (i.e., a single image containing a collage of key video frames) for all the collected videos using the method described in [16]. Storyboards enabled us to quickly determine whether a video contained substantial surgical footage, thereby avoiding the need for complex analysis on the entire sequence.
@@
[…]
Table 1: Performance of the video storyboard classification models across five folds.
@@ B.2 Video Selection and Trimming
[…]
B.3 Video Preprocessing
We use our trained frame classifier to detect and remove the intraoperative non-surgical frames. Additionally, we manually annotated 2719 surgical frames with 4584 non-surgical bounding-box instances and trained a YOLOv8 Nano model [31] to detect and obliterate the non-surgical content in surgical frames. After this process, we manually quality control all the remaining curated videos.
[…]
Your expertise includes identifying and categorizing surgical interventions based on clinical descriptions and procedural contexts.
Here is a list of 35 possible surgical procedure types: pancreatectomy, pancreaticoduodenectomy, splenectomy, ampullectomy, hepatectomy, nephrectomy, low anterior resection, colectomy, abdominoperineal resection, pulmonary lobectomy, hartmanns, prostatectomy, gastric bypass, duodenal switch, gastrectomy, small bowel resection, hernia repair, ulcer repair, cholecystectomy, appendectomy, ileocolic resection, cecectomy, myomectomy, hysterectomy, nissen fundoplication, adrenalectomy, thymectomy, rectopexy, adhesiolysis, esophagectomy, cystectomy, jejunostomy, ileorectal anastomosis, kidney transplant, vaginectomy.
Based on the description of the surgical video: <video title>, determine the most likely procedure type from the list. Focus on matching the description to the procedure type that best aligns with the terminology and context provided.
Figure 2: ChatGPT prompt employed to match video titles to procedure types. The title of the video to be matched is inserted where the $<$video title$>$ tag is located.
B.4 Video Annotation
For the surgery type, a video is considered to be robotic if the video title includes any of the following keywords: Robotic, Robot, Robo, Hugo, Versius, Senhance, Telerobotic, Console, and da Vinci. The search for these terms in the video titles was case-insensitive. The remaining videos, with titles that do not include any of these keywords, are manually verified to ensure that they are manual surgical procedures. For the surgical procedure type, we cross-reference the video titles with the predefined list of procedures. When exact title matches are not found, we leverage the capabilities of the ChatGPTv4 API to perform a more nuanced analysis, incorporating a customized prompt as shown in Fig. 2. All annotations are manually quality controlled.
[…]
Table 2: Performance of frame classification models across five folds.
@@ B.4 Video Annotation
Appendix C LemonFM Pretraining Details
This section provides an overview of the pretraining process for the LemonFM model.
LemonFM pretraining. For LemonFM pretraining, we ran on an Ubuntu 22.04.5 LTS node with eight NVIDIA V100 GPUs (32GB each), using a batch size of 24 per GPU under PyTorch 2.5.1+cu124 (CUDA 12.4). We employed AdamW as the optimizer with a teacher temperature of 0.04, fp16 precision, an initial learning rate of 5e-4 (after warm-up), a minimum learning rate of 1e-6, and a random seed of 30. The model was trained for 60 epochs, including 10 warm-up epochs. This number of epochs enables the model to converge and stabilize its training loss on our dataset and pretext task. The model with the lowest training loss was selected as the final model.
Appendix D Downstream Task Details
In this section, we provide details on the datasets used for evaluating each downstream task, including their respective data splits, as well as training configurations for linear probing and full fine-tuning settings. The experiments were conducted on an Ubuntu 22.04.5 LTS node equipped with an Intel IceLake Xeon CPU (72 vCPUs) and two NVIDIA A100 GPUs (each with 80 GB of memory), using PyTorch 2.5.1+cu124 (CUDA 12.4).
D.1 Data splits
Surgical phase recognition. For AutoLaparo, we followed the standard split of ten training videos, four for validation, and seven for testing [52]. For M2CAI16, we followed the data splits in [26, 51, 30], dividing the dataset into 27 training videos and 14 testing videos, respectively. For Cholec80, we adopted the data splits in [51, 12, 35, 27], allocating 40 videos for training, eight for validation, and 32 for testing.
Surgical tool presence detection. For GraSP, we adopted the data splits specified in [4], with four videos for training, four for validation, and five for testing. The data split for Cholec80 is identical to those employed in the surgical phase recognition task.
Surgical action recognition. For CholecT50, we followed the formal data split as proposed in [43].
Surgical semantic segmentation. For CholecSeg8k, we followed previous works [25, 18] and use 75% of the videos for training and 25% of the videos for testing (videos 12, 20, 48 and 55).
D.2 Linear probing
We used the teacher model from LemonFM as our backbone. The images were resized to $\times 224$ 224 for all downstream tasks. The downstream models with linear heads were trained with a batch size of 512, an initial learning rate (LR) of 1e-3, the AdamW optimizer, random seed of 30, and cross-entropy loss. An early stopping criterion was applied, stopping training after 10 epochs if the validation loss shows no improvement.
Table 3: Performance of bounding box detection models across five folds.
Fold Images Instances Precision Recall mAP50 mAP50-95
@@ D.2 Linear probing
D.3 Full fine-tuning